# SMILES has to go : Representation of Molecules via Algebraic Data Types

### v1

@@ 6 Conclusion
In this work, we have presented a novel approach to molecular representation using algebraic data types (ADTs) embedded in a pure, type-safe, functional programming language like Haskell. This ADT-based representation offers several compelling advantages over existing string-based representations such as SMILES and SELFIES. Firstly, the ADT representation inherently defines a grammar for molecular structures, ensuring that only semantically meaningful and valid molecules can be constructed. This eliminates the need for additional validation steps and reduces the likelihood of generating nonsensical or impossible structures. Secondly, the ADT representation is highly extensible, allowing for the seamless incorporation of additional molecular features and properties without requiring the creation of entirely new syntaxes or grammars. This flexibility enables the representation to evolve and adapt to the ever-expanding knowledge and requirements of cheminformatics. Thirdly, the ADT representation naturally supports the inclusion of 3D atomic coordinate data, enabling the calculation of crucial geometric properties such as dihedral angles and bond lengths. This is particularly advantageous for machine learning tasks that rely on 3D structural information, as opposed to SMILES and SELFIES which primarily focus on 2D connectivity. Furthermore, the ADT representation allows for the accurate modeling of resonance structures, tautomers, and organometallic compounds by incorporating the concept of delocalized electrons and multi-center bonding, inspired by the Dietz representation. This level of generality surpasses the capabilities of SMILES and SELFIES, which struggle to represent such complex molecular features in a single, unified structure. Additionally, the integration of the ADT representation with probabilistic programming paradigms opens up new possibilities for performing Bayesian inference and uncertainty quantification over molecular structures. By combining the expressive power of ADTs with the ability to represent and reason about probability distributions, we can develop rich, probabilistic models of molecules that capture the inherent uncertainty and variability associated with molecular properties. This probabilistic treatment is particularly valuable for drug discovery and design, where the ability to quantify uncertainty and incorporate prior knowledge is crucial for making informed decisions. Moreover, the use of a lazy functional programming language like Haskell offers potential efficiency improvements by leveraging lazy evaluation strategies. This is especially beneficial when performing simulations or sampling from molecular orbitals, as computations are only performed when strictly necessary. Finally, the ADT representation, being embedded in a statically-typed language, benefits from strong type safety guarantees. This helps catch potential errors at compile-time, reducing the chances of runtime failures and improving the overall reliability and maintainability of cheminformatics software. In conclusion, the proposed ADT-based molecular representation offers a powerful and flexible alternative to existing string-based representations like SMILES and SELFIES. By leveraging the strengths of functional programming, type safety, and extensibility, this approach has the potential to revolutionize the way we represent, manipulate, and reason about molecular structures in the field of cheminformatics.
7 Data, Funding, Acknowledgements and Abbreviations
@@ 7.1 Availability of data and materials
The library with instructions for its basic usage can be found at the following URL: https://github.com/oliverjgoldstein/chemalgprog
7.2 Funding
UK EPSRC DTP Scholarship.
@@ 7.3 Competing interests
There are no competing interests.
7.4 Authors Contributions
Oliver Goldstein provided the impetus and concept, as well as the entirety of the implementation.
7.5 Acknowledgements
I am grateful for conversations with Michael Harding about chemistry, and Sam March for helping with the restructuring and re-writing of parts of the article. I am grateful to Jack Liell-Cock for his helpful discussions about Haskell and graphs. Furthermore, I am grateful for the email exchanges with Dr. Richard Apodeca and his blog.
7.6 Author’s information
Oliver Goldstein is a Research Student at Oxford University reachable at oliverjgoldstein@gmail.com or oliver.goldstein@reuben.ox.ac.uk.
7.7 List of abbreviations
ADT(s) = Algebraic Data Type(s)
SMILES = Simplied Molecular Input Line Entry System

### v2

@@ 8 Conclusion
In this work, we presented a novel approach to molecular representation using algebraic data types (ADTs) embedded in the purely functional, type-safe, programming language Haskell.
We discussed contemporary representations, focusing primarily on the string based representations SMILES and SELFIES, and their short-comings for the use in machine learning.
By implementing the ADT using the Dietz representation for constitution of molecules, and atomic coordinate data for 3D information, the ADT greatly expands the chemical space which can be represented, and hence explored in cheminformatics tasks when compared to existing representations. In addition, the ADT allows the optional inclusion of quantum mechanical information, such as orbitals and electronic configurations, providing a level of detail unmatched in current representations.
As a general model for molecular representation, the ADT can represent complex, multi-centre, electron-deficient and delocalised bonding, as well as advanced stereoechemical features, and “zero order” bonds such as ionic and hydrogen bonds, without requiring extensive labeling for new bond types or oxidation states.
Through Haskell’s type classes and functional programming paradigm, we provide a rich and robust semantic context for molecular structure, and an efficient framework for computational tasks. Molecular data adheres to the constraints of the molecular representation, preventing nonsensical molecules from being formed, and going beyond the limitations of strings as a representation.
To demonstrate the ADT and how it can be extended, we implemented reaction networks and explore group based properties of molecular symmetry. We also showed integration of the ADT with probabilistic programming, showing that the ADT can be appropriately used for Bayesian inference, and that machine learning processes can operate directly on the program or grammar itself, without the need for external context or to encode the representation to other formats.
The ADT presented was not intended to be a definitive representation, but a valuable representational concept for cheminformaticians. Nevertheless the ADT offers a powerful and flexible alternative to existing string-based representations like SMILES and SELFIES, for the representation of, manipulation of, and reasoning about molecular structures in cheminformatics.
9 Declarations
@@ 9.1 Availability of data and materials
All code used to implement the Algebraic Data Type and the experiments in this paper is available via github, at https://github.com/oliverjgoldstein/chemalgprog